For given functions f(x) = 18 - 3x, and g(x) + 18 - 2xsquared
A. What is f(-6)? B. What is g(-3)?

Answer by KnightOwlTutor(293)

Plug in the value and solve.

What is f(-6)? f(x) = 18 - 3x =18+18=36 Remember (-)(-)=(+)
What is g(-3)? g(x) = 18 - 2xsquared =18-2(-3)(-3)=18-18=0